OpenStudy (anonymous):

you spin a spinner 48 times.out of the 48 trials, the pointer lands 6 times on section 2.what is the experimental probability of the pointer landing on section 2?

OpenStudy (anonymous):

6/48 = 1/8 = 12.5%